OpenStudy (anonymous):

What is the square root of (y/125)? Please show work.

OpenStudy (vijay):

\[1/5 * \sqrt{y/5}\]

OpenStudy (anonymous):

sqrt(y)/sqrt(125) = sqrt(y)/[sqrt(25)sqrt(5)] = sqrt(y) / (5sqrt(5)) = sqrt(5y) / (5*5) = [sqrt (5y)] /25

OpenStudy (anonymous):

y^1/2/5*5^1/2